# Time Distributed Deep Learning models for Purely Exogenous Forecasting. Application to Water Table Depth Prediction using Weather Image Time Series

### v1

@@ 2.3.1 TDC module
Along with the convolutional operations, the one hot encoding (OHE) of the corresponding month of $\tau$ is computed using 11-dimensional vector [^9] Encodes 12 exclusive categories into 12-dimensional binary vector would have brought a problem of linear dependence, i.e. perfect collinearity.; this lets the network to take account of seasonality behaviours.
The output of the TDC module is then, $\forall\tau\in[t-1,t-T]$, the concatenation of the OHE and the max pooling output. This is a multivariate time series with D variables, here named Time Distributed Hidden Representation. In our case study, $D$ is equal to 27 (16 max pooling dimension plus 11 OHE); and $P$ is equal to 3, i.e. the three weather variables employed: total precipitation, maximum temperature, and minimum temperature. The TDC module is employed in both two models, TDC-LSTM and TDC-UnPWaveNet, with the same structure and hyperparameters.
Figure 7: TDC module.
@@ 2.3.2 TDC-LSTM model
The TDC-LSTM model uses the TDC module and then a Sequential Module as depicted in Figure 8. In detail, a first bottleneck layer made of a TD fully connected followed by a Leaky-ReLU activation reduces the channel dimension to 16. This decreases the number of parameters in the subsequent layers, and thus it helps also in mitigating overfitting [31, 43]. Then, a 1D spatial dropout with probability $0.15$ is employed as a regularization technique. Instead of the classical dropout, the spatial dropout zero-out entire channels and not single element inside channels; this is done for facing correlation issues between consecutive elements in a channel [80]. A single LSTM layer with 32 units is then adopted to model the temporal relations, leaky-ReLU and sigmoid are employed as activation functions for the gates inside the LSTM cells. Leaky-ReLU has proved to be more effective than tanh in this task providing better results. A fully connected layer with 8 neurons and leaky-ReLU is then used to reduce the dimensionality of the 32-dimensional output of the LSTM. The last output layer computes an affine transformation and outputs the water table depth. The TDC-LSTM model has in total 9705 parameters, 6272 of which are from the LSTM layer.
@@ 2.3.3 TDC-UnPWaveNet model
The structure of the TDC-UnPWaveNet is very similar to the TDC-LSTM. The TDC module is still the same, however, the Sequential Module adopts the UnPWaveNet for learning temporal relations. Figure 6(b) depicts the architecture of the TDC-UnPWaveNet models. As for the TDC-LSTM, the first layer is still a bottleneck layer which reduces the channel dimension to 16. The bottleneck layer here is implemented as a 1x1 convolution, but as already stated in Section 2.2, it is equivalent to a TD fully-connected layer with the number of neurons equal to the number of filters of the 1x1 convolution. Then, the spatial dropout with probability $0.15$ is applied and its output is fed into the UnPWaveNet module (Figure 6(b)) which outputs an 8-dimensional vector that is fed into the last output layer identical to the TDC-LSTM model. The UnPWaveNet module is implemented using 5 layers of unpadded dilated convolution with 32 filters of size 4, and 1x1 convolution with 8 filters. The dilation is set $2^{l-1}$ for each layer $l$; in this way, the last (5th) layer has a dilation of 16. The TDC-UnPWaveNet model has in total 17915 parameters, 14746 of which are from the UnPWaveNet layer.
@@ 2.4.1 Preprocessing
The raw weather raster images covered the entire Piemonte region, to focus on the catchment all the images were clipped on the ROI maintaining a squared shape which is easier to handle with CNN. A box with a lower-left corner in coordinate (6.90°E;44.35°N) and higher-right in (7.79°E;44.84°N) was adopted to clip the images. Concerning the temporal resolution, we set a weekly time step for the predictions, and then both the target and features were aggregated computing weekly averages.
In a time series task in which lagged features are employed, inserting gaps between the training, validation and test sets is common to prevent data leakage and performance overestimation. For example, in the case of an autoregressive model which uses lags up to $t-T$ as features to predict the target at time $t$, it is usual to discard the $T$ time steps between sets. In the present case study, even if the models do not use autoregressive terms but only exogenous (weather) variables, we think that a gap is still needed because the last observations in a set use almost the same predictor as the first observation in the next set. Then, especially in cases where $T$ is large, this affects the independence of consecutive sets, and then, the unbiasedness of the performance metrics. More formally it is possible to speak about leakage from training examples as described in [81]. Many related studies as [25, 27] did not mention this issue, and built consecutive and overlapping sets. Here, instead, in a more precautionary fashion, a gap of $T$ time step is considered for defining sets. In other words, we have constrained that features used in a set can not be used as predictors also in a subsequent set.
@@ 2.4.2 Hyperparameters and training
3 Results
Many performance metrics have been computed following Equations 2 3 4 5 6 7 89. In these equations $\overline{y}$, $y_{min}$, $y_{max}$ represent respectively the target mean, minimum and maximum computed over the training set. NSE and KGE have been frequently adopted in hydrological modeling studies. More in detail, NSE shows how a model performs with respect to a naive estimator, which is usually the mean ($\overline{y}$). The NSE has an intrinsic benchmark set at 0, that is when the model performs as well as the naive estimator. The KGE [82] is a different concept, it is the the euclidean distance between the vector defined by metrics $\rho$ (as in Equation 7),$\alpha$ and, $\beta$ and the vector with the best achievable metrics ($\rho=1$, $\alpha=1$ and, $\beta=1$). In practice, the KGE is a measure which takes into account more aspect of the prediction, i.e. the Pearson correlation, the bias and the variance. The KGE has no intrinsic benchmark as the NSE and, as pointed out in [83], these two metrics are not directly comparable. In [83] authors stated that if the benchmark is set as the NSE. i.e. a fixed mean estimation, then the cut-off point of the KGE is $-0.41$, after which the model is performing better than the naive estimator. For both NSE and KGE the maximum value is 1 and the higher the values are the better a model is performing.
Table 3 shows all the performance metrics computed over the test sets for the TDC-LSTM and TDC-UnPWaveNet models, and Figures 11(a) 11(b) 11(c) show the ensemble mean prediction. It is difficult to define a clear winner between the TDC-LSTM and TDC-UnPWaveNet, also because it seems that the two models have captured different aspects of the phenomenon to be modelled. In the case of Vottignasco and Racconigi sensors, the TDC-LSTM has performed better in terms of RMSE, BIAS, MAPE and NSE, however, the TDC-UnPWaveNet has been better for correlation and KGE. The TDC-UnPWaveNet has appeared to be more able to predict the actual temporal evolution of the ground truth, while the TDC-LSTM has been better in terms of the biasedness of predictions. This is very clear in Figure 11(c) in which the TDC-UnPWaveNet follow accurately the temporal evolution ($\rho=0.95$) of the ground truth but a bias is clearly visible. Another example could be the drop in the Vottignasco series (Figure 11(a)) around 2022-10-01. Here, the TDC-UnPWaveNet has predicted correctly a more prolonged drop and more in line with the actual values, instead the TDC-LSTM has predicted a very accurate decrease of the depth, however, it incorrectly has predicted the recovery too early, probably driven by training-related memory (i.e. overfitting). In this terms, it seems that the TDC-UnPWaveNet could generalize better, at the cost of higher bias. For the Savigliano the TDC-UnPWaveNet has won, differences in $\rho$ and KGE are almost negligible and probably not significant from a statistical point of view. From Figure 11(b) the TDC-UnPWaveNet predictions show temporal dynamics more in line with the actual values, while the TDC-LSTM predictions appear to be far more erratic than the ground truth. Notwithstanding, for that series, the differences are very narrow and both the models have performed very well.
Figure 11: a) Ensemble mean predictions on the test set of the Vottignasco sensor. b)Ensemble mean predictions on the test set of Savigliano sensor c) Ensemble mean predictions on the test set of the Racconigi sensor. For all the plots the shadows represent two times the standard deviation of the ensemble predictions. For shortening the legends, LSTM stands for the TDC-LSTM model, while UnPWaveNet for the TDC-UnPWaveNet model. For ease the interpretation of the water table depth y axes has been reversed.
@@ 3 Results
Table 3: Performance metrics on test sets - Ensemble Mean Predictions.
@@ 4 Discussion
Concerning the test period, it has to be stressed out that 2022 was a very particular year in terms of weather conditions. In fact, our ROI suffered from a severe drought in the summer which lasted until the autumn of 2022. Thus predicting the water table depth, especially in Vottignasco and Racconigi, has been a very difficult task for our models, which have to predict an uncommon drop. This has been even more difficult given the absence of an explicit autoregressive term, which would have helped the model in anchoring to the most recent actual water table depth values.
As stated in Section 1.2 the choice of not using autoregressive terms in our proposed models has been guided only by a practical fact, i.e. water table depth data are updated on a semester basis. Then, using an autoregressive term would have made our model unusable in a practical scenario to predict the next week’s depth value. Furthermore, no anthropogenic pressure proxy (e.g. human water consumption) has been fed to our model, and this is because of the lack of such data in our ROI. Even if our models have yielded satisfactory results, they could be enhanced by the introduction of these additional inputs (as in [65, 16, 23, 25]), especially to improve the performance in anomalous scenarios like the summer and autumn of 2022. In fact, predicting such a drop could be very difficult looking only at the weather variable. For example, if the precipitations are scarce, likely, the anthropogenic pressure on the groundwater resource increases making groundwater resources decrease even more.
In [25] authors found the LSTM-based model more robust against initialization effects than CNN. We have found soft evidence of this. In fact, in our case study, this could be true for Vottignasco and Racconigi series, in which the ensemble standard deviation of the TDC-LSTM models (shadows in Figure 11(a) and 11(c)) seem to be lower than the TDC-UnPWaveNet ones. However, this is not true for the Savigliano test predictions, in which the TDC-UnPWaveNet ensemble standard deviation appears to be lower. Furthermore, in our case study, the more variability related to the initialization effect could be also caused by the higher number of parameters and the deeper architecture of the TDC-UnPWaveNet.
@@
What emerges from the present study is that both the TDC-UnPWaveNet and TDC-LSTM have produced satisfactory predictions but with different modeling abilities. The ability of the TDC-UnPWaveNet in modeling better the actual temporal dynamics could be due to the more complex structure and the higher number of parameters, which in the framework of DL remain negligible for both models. Furthermore, it should be considered that the TDC-UnPWaveNet is a CNN-based model, so it is highly parallelizable and time-efficient in terms of computations [39].
@@ 5 Conclusion
We have proposed two different DL models for predicting, in a many-to-one fashion, the water table depth of three sensors located in the Grana-Maira catchment (Piemonte, IT) from weather image time series. These models are made of two modules: a first Time Distributed CNN (TDC) and a Sequential Module. The TDC is the same for the two proposed models, and it extracts a vectorial representation (Time Distributed Hidden Representation) of the input image time series, i.e. it encodes each image available at each time step into a vector forming a hidden multivariate time series. For the TD-LSTM model, the Sequential Module is based on a classical LSTM layer; instead for the TD-UnPWaveNet model the sequential model is based on a new version of the WaveNet adapted here to output a series completely in the future and shorter than the input one - actually a many-to-one scenario for this case study.
In developing the UnPWaveNet, and facing the issue of different sequence lengths inside the architecture, we have designed a new Channel Distributed (CD) layer. The CD layer applies the same transformations to each channel individually (i.e. a translation of the concept of Time Distributed layer to channels). In this way, a sequence with many channels could be transformed into a sequence of a different length maintaining the channel-wise dimension. The CD layer, implemented in the UnPWaveNet with a fully connected cell, has proved to be efficient and effective: it has enabled to achieve very satisfactory results limiting the total number of parameters.

### v2

@@ 2.3.1 TDC module
Along with the convolutional operations, the one hot encoding (OHE) of the corresponding month of $\tau$ is computed using 11-dimensional vector [^7] Encoding 12 exclusive categories into a 12-dimensional binary vector would have brought to linear dependent features, i.e. perfect collinearity.; this lets the network to take account of seasonality behaviors.
The output of the TDC module is then, $\forall\tau\in[t-1,t-T]$, the concatenation of the OHE, and the max pooling output. This is a multivariate time series with D variables, here named Time Distributed Hidden Representation. In our case study, $D$ is equal to 27 (16 max pooling dimension plus 11 OHE); and $P$ is equal to 3, i.e. the three weather variables employed: total precipitation, maximum temperature, and minimum temperature. The TDC module is employed in both two models, TDC-LSTM and TDC-UnPWaveNet, with the same structure and hyperparameters.
Figure 7: Time Distributed CNN (TDC) module.
@@ 2.3.2 TDC-LSTM model
The TDC-LSTM model uses the TDC module and then a Sequential Module as depicted in Figure 8. In detail, a first bottleneck layer made of a TD fully connected followed by a Leaky-ReLU activation reduces the channel dimension to 16. This decreases the number of parameters in the subsequent layers, and thus it helps also in mitigating overfitting [37, 49]. Then, a 1D spatial dropout with probability $0.15$ is employed as a regularization technique. Instead of the classical dropout, the spatial dropout zero-out entire channels and not single element inside channels; this is done to face correlation issues between consecutive elements in a channel [86]. A single LSTM layer with 32 units is then adopted to model the temporal relations, leaky-ReLU, and sigmoid are employed as activation functions for the gates inside the LSTM cells. Leaky-ReLU has proved to be more effective than tanh in this task providing better results. A fully connected layer with 8 neurons and leaky-ReLU is then used to reduce the dimensionality of the 32-dimensional output of the LSTM. The last output layer computes an affine transformation and outputs the water table depth. The TDC-LSTM model has 9705 parameters in total.
@@ 2.3.3 TDC-UnPWaveNet model
The structure of the TDC-UnPWaveNet is very similar to the TDC-LSTM. The TDC module is still the same, however, the Sequential Module adopts the UnPWaveNet for learning temporal relations. Figure 9 depicts the architecture of the TDC-UnPWaveNet models. As for the TDC-LSTM, the first layer is still a bottleneck layer which reduces the channel dimension to 16. The bottleneck layer here is implemented as a 1x1 convolution, but as already stated in Section 2.2, it is equivalent to a TD fully-connected layer with the number of neurons equal to the number of filters of the 1x1 convolution. Then, the spatial dropout with probability $0.15$ is applied, and its output is fed into the UnPWaveNet module (Figure 6(b)) which outputs an 8-dimensional vector that is fed into the last output layer identical to the TDC-LSTM model. The UnPWaveNet module is implemented using 5 layers of unpadded dilated convolution with 32 filters of size 4, and 1x1 convolution with 8 filters. The dilation is set $2^{l-1}$ for each layer $l$; in this way, the last (5th) layer has a dilation of 16. The TDC-UnPWaveNet model has 17915 parameters in total.
@@ 2.4.1 Preprocessing
The raw weather raster images covered the entire Piedmont region, to focus on the catchment all the images were clipped on the ROI maintaining a squared shape which is easier to handle with CNNs. A box with a lower-left corner in coordinate (6.90°E;44.35°N) and higher-right in (7.79°E;44.84°N) was adopted to clip the images. Concerning the temporal resolution, we set a weekly time step for the predictions, and then both the target and features were aggregated computing weekly averages.
In a time series task in which lagged features are employed, inserting gaps between the training, validation, and test sets is common to prevent data leakage and performance overestimation. For example, in the case of an autoregressive model that uses lags up to $t-T$ as features to predict the target at time $t$, it is usual to discard the $T$ time steps between sets. In the present case study, even if the models do not use autoregressive terms but only exogenous (weather) variables, we think that a gap is still needed because the last observations in a set use almost the same predictor as the first observation in the next set. Then, especially in cases where $T$ is large, this affects the independence of consecutive sets, and then, the unbiasedness of the performance metrics. More formally it is possible to speak about leakage from training examples as described in [87]. Many related studies as [31, 33] did not mention this issue, and built consecutive and overlapping sets. Here, instead, in a more precautionary fashion, a gap of $T$ time step is considered for defining sets. In other words, we have constrained that features used in a set can not be used as predictors also in a subsequent set.
@@ 2.4.2 Hyperparameters and training
2.4.3 Evaluation metrics
Many performance metrics have been computed following Equations 2345678, and 9. In these equations $\overline{y}$, $y_{min}$, $y_{max}$ represent respectively the target mean, minimum, and maximum computed over the training set. NSE and KGE have been frequently adopted in hydrological modeling studies. More in detail, NSE shows how a model performs with respect to a naive estimator, which is usually the mean ($\overline{y}$). The NSE has an intrinsic benchmark set at 0, which is when the model performs as well as the naive estimator. The KGE [88] is a different concept, it is the Euclidean distance between the vector defined by metrics $\rho$ (as in Equation 7),$\alpha$ and, $\beta$ and the vector with the best achievable metrics ($\rho=1$, $\alpha=1$ and, $\beta=1$). In practice, the KGE is a measure that takes into account more aspects of the prediction, i.e. the Pearson correlation, the bias, and the variance. The KGE has no intrinsic benchmark as the NSE and, as pointed out in [89], these two metrics are not directly comparable. In [89] authors stated that if the benchmark is set as the NSE. i.e. a fixed mean estimation, then the cut-off point of the KGE is $-0.41$, after which the model performs better than the naive estimator. For both NSE and KGE the maximum value is 1 and the higher the values the better a model is performing.
@@ 2.4.3 Evaluation metrics
3 Results
Table 3: Performance metrics on test sets computed on the ensemble mean predictions. The last row reports the means and the standard deviations in parenthesis over the three sensors.
@@ 3 Results
Mean ($\sigma$) TDC-LSTM 0.27 0.08 -0.18 -0.05 0.04 0.93 0.91 0.41
(0.12) (0.01) (0.05) (0.01) (0.01) (0.03) (0.02) (0.07)
TDC-UnPWaveNet 0.31 0.08 -0.25 -0.05 0.05 0.96 0.90 0.43
(0.19) (0.03) (0.19) (0.04) (0.03) (0.01) (0.08) (0.05)
Table 3 shows the performance metrics computed over the test sets for the TDC-LSTM and TDC-UnPWaveNet models, the bottom row reports the mean performances over the three sensors with the corresponding standard deviations $\sigma$. Figures 11(a) 11(b) 11(c) show the ensemble mean prediction. It is difficult to define a clear winner between the TDC-LSTM and TDC-UnPWaveNet, also because it seems that the two models have captured different aspects of the phenomenon to be modeled. In the case of Vottignasco and Racconigi sensors, the TDC-LSTM has performed better in terms of RMSE, BIAS, MAPE, and NSE, however, the TDC-UnPWaveNet has been better for correlation and KGE. The TDC-UnPWaveNet has appeared to be more able to predict the actual temporal evolution of the ground truth, while the TDC-LSTM has been better in terms of the biasedness of predictions. This is very clear in Figure 11(c) in which the TDC-UnPWaveNet follows accurately the temporal evolution ($\rho=0.95$) of the ground truth but a bias is clearly visible. Another example could be the drop in the Vottignasco series (Figure 11(a)) around 2022-10-01. Here, the TDC-UnPWaveNet has predicted correctly a more prolonged drop and more in line with the actual values, instead the TDC-LSTM has predicted a very accurate decrease of the depth, however, it incorrectly has predicted the recovery too early, probably driven by training-related memory (i.e. overfitting). In this terms, it seems that the TDC-UnPWaveNet could generalize better, at the cost of higher bias. For the Savigliano the TDC-UnPWaveNet has won, differences in $\rho$ and KGE are almost negligible and probably not significant from a statistical point of view. From Figure 11(b) the TDC-UnPWaveNet predictions show temporal dynamics more in line with the actual values, while the TDC-LSTM predictions appear to be far more erratic than the ground truth. Notwithstanding, for that series, the differences are very narrow and both the models have performed very well.
Figure 11: Ensemble mean predictions on the test set of sensors in a) Vottignasco, b) Savigliano, c) Racconigi. For all the plots the shadows represent two times the standard deviation of the ensemble predictions. For shortening the legends, LSTM stands for the TDC-LSTM model, while UnPWaveNet for the TDC-UnPWaveNet model. For ease the interpretation of the water table depth y axes has been reversed.
@@ 4 Discussion
Concerning the test period, it has to be highlighted that 2022 was a very particular year in terms of weather conditions. Indeed, our ROI suffered from a severe drought in the summer which lasted until the autumn of 2022. Thus predicting the water table depth, especially in Vottignasco and Racconigi, has been a very difficult task for our models, which have to predict an uncommon drop. This has been even more difficult given the absence of an explicit autoregressive term, which would have helped the model in anchoring to the most recent actual water table depth values.
As stated in Section 1.2 the choice of not using autoregressive terms in our proposed models has been guided only by a practical fact, i.e. water table depth data are updated on a semester basis. Then, using an autoregressive term would have made our model unusable in a practical scenario to predict the next week’s depth value. Furthermore, no anthropogenic pressure proxy (e.g. human water consumption) has been fed to our model, and this is because of the lack of such data in our ROI. Even if our models have yielded satisfactory results, they could be enhanced by the introduction of these additional inputs (as in [71, 90, 29, 31]), especially to improve the performance in anomalous scenarios like the summer and autumn of 2022. Indeed, predicting such a drop could be very difficult looking only at the weather variable. For example, if the precipitations are scarce, likely, the anthropogenic pressure on the groundwater resource increases making groundwater resources decrease even more.
In [31] authors found the LSTM-based model more robust against initialization effects than CNN. We have found soft evidence of this. Indeed, in our case study, this could be true for Vottignasco and Racconigi series, in which the ensemble standard deviation of the TDC-LSTM models (shadows in Figure 11(a) and 11(c)) seem to be lower than the TDC-UnPWaveNet ones. However, this is not true for the Savigliano test predictions, in which the TDC-UnPWaveNet ensemble standard deviation appears to be lower. Furthermore, in our case study, the more variability related to the initialization effect could be also caused by the higher number of parameters and the deeper architecture of the TDC-UnPWaveNet.
@@
What emerges from the present study is that both the TDC-UnPWaveNet and TDC-LSTM have produced satisfactory predictions but with different modeling abilities. The ability of the TDC-UnPWaveNet to model better the actual temporal dynamics could be due to the more complex structure and the higher number of parameters, which in the framework of DL remain negligible for both models. Furthermore, it should be considered that the TDC-UnPWaveNet is a CNN-based model, so computationally more efficient thanks to the possibility of parallelization [45].
@@ 5 Conclusion
We have proposed two different DL models for predicting, in a many-to-one fashion, the water table depth of three sensors located in the Grana-Maira catchment (Piedmont, IT) from weather image time series. These models are made of two modules: a first Time Distributed CNN (TDC) and a Sequential Module. The TDC is the same for the two proposed models, and it extracts a vectorial representation (Time Distributed Hidden Representation) of the input image time series, i.e. it encodes each image available at each time step into a vector forming a hidden multivariate time series. For the TD-LSTM model, the Sequential Module is based on a classical LSTM layer; instead for the TD-UnPWaveNet model the sequential model is based on a new version of the WaveNet adapted here to output a series completely in the future and shorter than the input one - actually a many-to-one scenario in this case study.
In developing the UnPWaveNet, and facing the issue of different sequence lengths inside the architecture, we have designed a new Channel Distributed (CD) layer. The CD layer applies the same transformations to each channel individually (i.e. a translation of the concept of Time Distributed layer to channels). In this way, a sequence with many channels could be transformed into a sequence of a different length maintaining the channel-wise dimension. The CD layer, implemented in the UnPWaveNet with a fully connected cell, has proved to be efficient and effective: it achieves very satisfactory results limiting the total number of parameters.
@@ 7 Software and data availability
8 CRediT
Metteo Salis: Conceptualization, Methodology, Software & Data curation, Formal Analysis, Investigation, Writing, Validation, Visualization; Abdourrahmane M. Atto: Conceptualization, Methodology, Supervision, Review, Validation; Stefano Ferraris: Conceptualization, Supervision, Review, Validation; Rosa Meo: Conceptualization, Methodology, Supervision, Review, Validation.